74.) Sketch the ellipse. Identify the center, vertices, and foci.
9(x+4)^2+(y+7)^2=81

Answer by lwsshak3(11628)

Sketch the ellipse. Identify the center, vertices, and foci.
9(x+4)^2+(y+7)^2=81
**
Standard form of equation for an ellipse with vertical major axis:
(x-h)^2/b^2+(y-k)^2/a^2=1,a>b, with(h,k) being the (x,y) coordinates of the center.
..
9(x+4)^2+(y+7)^2=81
divide by 81
(x+4)^2/9+(y+7)^2/81=1
center: (-4,-7)
a^2=81
a=9
vertices: (-4,-7±a)=(-4,-7±9)=(-4,2) and (-4,-16)
..
b^2=9
b=3
..
c^2=a^2-b^2=81-9=72
c=√72≈8.5
Foci: (-4,-7±c)=(-4,-7±8.5)=(-4,1.5) and (-4,-15.5)
..
see graph below as a visual check:
y=±(81-9(x+4)^2)^.5-7